TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.MariaPazCasasLino on 2007-05-31 - 18:34

Powering test in sector 7-8. 31 May - 01 June 2007

COMPLETED ON 31/05/07

Q5 Complete cycle of Q5 up to nominal current (4310 A) PNO.8d. Test OK.
Correctors of Q4 and Q5
Battery test of six correctors: PNO.3. Test OK.
Battery test of six correctors: PNO.4 (SPA). Not OK for FGC software problem.
Q4 corrector released by MPP after quench analysis. PNO.3. Test OK.
D2, Q4, Q5 Powering to nominal cycle of all the matching section ML8 together. Test not OK for quench in D2 and Q4. The quench occurred when all the circuits were at about 80% of the nominal current. Maximum was 5268 A for D2.
ElQA 13kA Completed for RQF and RQD.
In RB discharged at 1.7 kV (1.9 kV nominal). Cryo instrumentation was disconnected and next discharged was at 300 V. In the evening it was foreseen to repeat it with all the electronic disconnected.

PLAN FOR 01/06/07

ElQA HV tests dipole circuit
RQ6.R7 High voltage test to be repeated
RQ6.R7 in DFBMH Once ElQA finished and insulation plug restored
* QPS controller to be reprogrammed
* Current leads to be connected
* PCC Circuit Config
Matching Section ML8 Blocked by MPP till deeper understanding of the quench event. No powering tests till MPP releases it.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-31 - 17:54

Beam Dump System - Installation & Commissioning - 22nd May

PDF RAT_20070522.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x05x31x17x53/RAT_20070522.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-31 - 17:52

Short Circuit Test in UA27 - 31st May

Microsoft Excel Spreadsheet SOC_II_8hr_heatrun_UA27.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x31x17x49/SOC_II_8hr_heatrun_UA27.xls

PDF RAT_20070531.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x31x17x49/RAT_20070531.pdf

Comments:

EE - IST    
13kA EE DONE 16-04-07
600A EE DONE 23-04-07
     
Circuits >2kA    
Powering tests PC 4-8kA DONE 11-05-07
8h run 4-8kA ONGOING 29-05-07
Interlock tests RQD&F ONGOING 30-05-07
Powering tests RQD&F ONGOING 30-05-07
Preparation RB.A23 DONE 19-04-07
Interlock tests RB.A23 DONE 25-05-07
Powering tests RB.A23 ONGOING 30-05-07
8h run >= 8kA - 04-06-07
     
Circuits < 2kA    
120A powering tests DONE 03-05-07
600ANOEE powering tests DONE 08-05-07
600A-EE interlock tests DONE 03-05-07
600A-EE powering tests DONE 18-05-07
8 hr run PC < 2kA DONE 29-05-07
     
24hr heat run test - 07-06-07
     
Polarity Check < 8KA - 30-05-07
AUG- UPS test - 08-06-07

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-30 - 18:52

Powering test in sector 78 - 30-31 May 2007

COMPLETED ON 30/05/07

CORRECTOR CIRCUITS (Q4,Q5) RCBYVS4.L8B1; RCBYVS4.L8B2; RCBYV4.L8B1 went through PNO.1, PNO.2 and partially PNO.3
RCBYVS4.L8B2 had a training quench during PNO.2
RCBCH5.L8B1; RCBCV5.L8B2; RCBYH4.L8B2; RCBYHS4.L8B1 went partially through PNO.3
RQ4.L8
Unbalanced ramp up to nominal and unbalanced quench (PNO.8c)

Third trial for the unbalanced cycle up to nominal: B1 and B2 got to nominal current. In the ramp down of B1, due to a too the fast negative slope the converter tripped. B2 slow power abort commanded by PIC.
(Since B2 reached nominal current, there is not need to repeat the quench at nominal current in both apertures, PNO.7a is considered OK)
Provoked unbalanced quench with (B1 at nominal). PNO.8c OK.
Current cycle up to nominal PNO.9
RQ5.L8 Provoked unbalanced quench with (B1 at nominal). Ramp down from nominal this time worked OK. PNO.8c OK.

PROPOSAL OF THE ACTIVITY OF THURSDAY 31/05/07

CORRECTOR CIRCUITS (Q4,Q5) RCBYVS4.L8B1; RCBYVS4.L8B2; RCBYV4.L8B1; RCBCH5.L8B1; RCBCV5.L8B2; RCBYH4.L8B2; RCBYHS4.L8B1; should go through PNO.3 and PNO.4
RQ4.L8
RQ5.L8
Current cycle up to nominal PNO.9 on Q5.
As soon as they are released, the squeeze function test will be performed on these 2 circuits.

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-29 - 18:52

Powering test of sector 78 - minutes of the meeting of 29th Mat 2007 afternoon

PDF RAT_20070529b.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x29x18x50/RAT_20070529b.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-29 - 17:27

Powering Tests - Sector 7-8 - 30/05/2007

Yesterday was devoted to perform tests on RD2.L8 and RQ4.L8

    • RD2.L8 has been successfully commissioned up to PNO.8d (Current cycle up to nominal current)
    • Missing tests for RD2.L8: Fast Power Abort from PIC (PNO.9). The detection threshold needs to be modified. Possibly done while powering at unison (like PNO.9 for RQ5)
    • RQ4.L8:
        • Unbalanced cycle up to nominal (3610 A) for B1 and 2000 A for B2. While ramping B1 converter tripped. The attempt of PO expert to recover control made the quench detector trip (same problem as last week)
        • Second trial of unbalanced cycle up to nominal: with B1 at nominal current and B2 ramping up to nominal, the latter quenched at 3590 A. Test failed.

Proposal for the activity of today (30/05/2007)

  • Complete P2N of all the 120-80A correctors of Q4 and Q5.
  • Continue P2N of RQ4.L8. Two options: * Unbalanced ramp up and quench with both apertures at nominal current (PNO.7a) or/and * Unbalanced cycle and quench with one aperture at nominal and the other at 50%.

  • ElQA activities: Morning: * Full qualification of the 3 circuits selected on DFBAN * TFM on circuits powered via DFBMH * preparation for the HV qualification of the 600A circuits Afternoon * HV qualification on 600 A circuits powered via DFBAO

  • QPS: Interlock tests .

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-29 - 11:12

Short Circuit Test UA27 - 29th May

PDF RAT_20070529.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x29x11x06/RAT_20070529.pdf

**********************************

8hr heat run ON GONIG: all circuits belonging to the powering sub-sector MR2 and A23 (but 13kA)

**********************************

EE - IST    
13kA EE DONE 16-04-07
600A EE DONE 23-04-07
     
Circuits >2kA    
Powering tests PC 4-8kA DONE 11-05-07
8h run 4-8kA ONGOING 29-05-07
Interlock tests RQD&F ONGOING 30-05-07
Powering tests RQD&F - 30-05-07
Preparation RB.A23 DONE 19-04-07
Interlock tests RB.A23 DONE 25-05-07
Powering tests RB.A23 - 30-05-07
8h run >= 8kA - 04-06-07
     
Circuits < 2kA    
120A powering tests DONE 03-05-07
600ANOEE powering tests DONE 08-05-07
600A-EE interlock tests DONE 03-05-07
600A-EE powering tests DONE 18-05-07
8 hr run PC < 2kA ONGOING 29-05-07
     
24hr heat run test - 07-06-07
     
Polarity Check < 8KA - 30-05-07
AUG- UPS test - 08-06-07

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2007-05-29 - 10:57

Powering Tests of Sector 7-8: minutes of the meetings of 29/05

PDF RAT_20070529.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x29x10x56/RAT_20070529.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-25 - 21:28

Powering Tests in Sector 7-8: activity of Friday 25/05.

The day was fully devoted to perform tests on D2.

The following tests have been executed on D2 (MANUAL execution, due to change in procedure):
  • PIC2 (SEQUENCER execution); test successful;
  • PLI1: ramp to 350 A with 10 A/s ramp rate, we stayed on the plateau 4 min and then provoked a FPA; test successful;
  • PLI2: ramp to 350 A with 10 A/s ramp rate, then ramp to 1000 A with 10 A/s, and quench driven by QPS, generating an artificial quench signal from one current lead detector; the test was successful;
  • PNO: during the first attempt to ramp to 3000 A, the power converter went to ‘faults’ and produced a FPA when reaching 1008 A; investigations of PO showed that a hardware limitation was present. The second attempt was successful:
    • Ramp to 350 A with 10 A/s ramp rate;
    • Continue ramping to 3000 A; plateau of 4 min;
    • Continue ramping to 6000 A (nominal current); plateau of 30 min to check voltage drift;
    • Provoked quench at nominal current.

Proposal for the activity of next week.

  • Continue powering Q4 and D2
  • Continue powering the correctors (4 circuits have already passed the first 2 steps, three more are available).
  • Possible powering of the whole matching section.
  • Powering of the arc 60 A orbit correctors, according to humidity conditions.

  • Continue installation of plastic bags on DFBAN and DFBMH.
  • Check the status of condensation of the 13kA leads of DFBAO which have copper laminations and no plastic bags; check the status of the 6kA lead of DFBMA which has copper laminations and no plastic bags.
  • Installation of ventilators on the 60 A orbit correctors leads.

  • ElQA on the selected circuits: see presentation of Davide in the last post.

  • Quench heaters qualification, without interference with ElQA.

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-25 - 21:15

Powering Tests of Sector 7-8: minutes of the meetings of 25/05.

PDF Q4.L8.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x25x21x07/Q4.L8.pdf

PDF RAT_20070525.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x25x21x07/RAT_20070525.pdf

PDF ELQA_CCC_25052007.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x25x21x07/ELQA_CCC_25052007.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-25 - 11:33

Short Circuit Tests in UA27 - 25th May

PDF RAT_20070525.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x25x11x30/RAT_20070525.pdf

EE - IST    
13kA EE DONE 16-04-07
600A EE DONE 23-04-07
     
Circuits >2kA    
Powering tests PC 4-8kA DONE 11-05-07
8h run 4-8kA - 29-05-07
Interlock tests RQD&F ONGOING 02-05-07
Powering tests RQD&F - 24-05-07
Preparation RB.A23 ON GOING 19-04-07
Interlock tests RB.A23 DONE 25-05-07
Powering tests RB.A23 - 30-05-07
8h run >= 8kA - 04-06-07
     
Circuits < 2kA    
120A powering tests DONE 03-05-07
600ANOEE powering tests DONE 08-05-07
600A-EE interlock tests DONE 03-05-07
600A-EE powering tests DONE 18-05-07
8 hr run PC < 2kA - 29-05-07
     
24hr heat run test - 07-06-07
     
Polarity Check < 8KA - 30-05-07
AUG- UPS test - 08-06-07

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-24 - 19:26

Powering tests in Sector 7-8: activity of Thursday 24/05.

The day was fully devoted to perform tests on Q4, which had experienced an unexpected quench at 300 A on Monday (after ramping at 600 A).

The following tests have been executed on Q4 (MANUAL execution, due to change in procedure):
  • ramp to 600 A with 5 A/s and quench driven by QPS, generating an artificial quench signal from one current lead detector; the test was successful;
  • ramp of B1 power converter to nominal current (3610 A) with 10 A/s ramp rate and simultaneous ramp of B2 power converter to 2000 A with nearly 5 A/s (unbalanced configuration):
    • when B1 current reached 1025 A, its reference current went anormally down to 0;
    • later analysis showed that a fault was sent from the converter and a consequent FPA was started;
    • PO attempt to get control of B2 power converter, to ramp it down contemporarily to 0, provoked a sudden voltage peak in B1, which initiated a quench;
  • at 11.48 (according to PIC supervision) a power cut occured on the general services of UA83, which eventually provoked the loss of controls on cryogenics and lead heating system;
  • after waiting for the good cryo conditions to be recovered, and in agreement with PO, QPS and MPP experts, we decided to ramp again Q4 with the same unbalanced cycle:
    • B1 at 3610 A with 10 A/s ramp rate and B2 to 2000 A with a nearly 5 A/s ramp rate;the test was successful and NO quench was activated;
    • while ramping B2 power converter to the same current as B1 (the idea was to stay on plateau for half an hour and then provoke a quench), and with a current in B2 of 2900 A, a quench generated; preliminary analysis indicated that it was started in B1, possibly due to electromagnetic coupling (this configuration has never been tested before).
  • end of the tests due to long recovery time.

Proposal for the activity of tomorrow.

  • Continue powering Q4 and D2.

  • Continue installation of plastic bags on 600 A current leads.
  • Lower the temperature on the 13 kA leads to check the absence of condensation with AT/MEL copper laminations.

  • ElQA on the selected circuits: list will be provided during the morning.

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-05-24 - 15:02

Powering Tests Sector 7-8 - Minutes 24/05/2007

PDF RAT_20070524b.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x24x14x54/RAT_20070524b.pdf

PowerPoint Q5_test_240507.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x24x14x54/Q5_test_240507.ppt

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-24 - 10:06

Sector 7-8 cooldown

After becoming the first magnet at 4.5K and in the kA range to be powered last week, the Q5 matching section quadrupole (MQM) left of Point 8 became the first one to reach nominal current on May 23 @ 15:04:28. The two apertures of the magnet were then powered at different current levels and a quench was provoked. Following that, the magnet was taken to nominal current. This magnet, which was suspected to require training quenches, reached nominal current after two quenches (3910A and 4224 A) before reaching nominal current (4310A) [see the HC Twiki pages for details].

Commissioning now proceeds with Q4 and then D2 as well as the associated orbit correctors the same region.

In the meantime, the cryo team has continued with the tuning of the pumping group and the 4.5K refrigerator and has started cooling down the arc down to 1.9K. This morning (Thursday, May 24) half of the arc was below lambda point. The Power Converter, Interlock, Current Lead, Electrical Quality Assurance and Quench Protection System teams have started the preparation for the powering of the arc early next week.

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-23 - 19:02

Powering Tests of Sector 7-8: activity of Wednesday 23/05.

Today the activity was concentrated on Q5, which was released by MPP after analysis of the quench of Monday morning.

The following tests have been executed on Q5 (MANUAL execution):
  • ramp to nominal (4310 A) with 5 A/s ramp rate: quenched occurred at 4210 A on B2 aperture; preliminary analysis confirmed it was a training quench;
  • ramp to nominal (4310 A) with 5 A/s ramp rate: NO quench; a plateau of 10 min was mantained for possible voltage drift recording;
  • keeping B1 power converter at nominal current, B2 power converter was lowered to 55% of the nominal current (2370 A); with this unbalanced configuration, a quench was driven by QPS generating an artificial quench signal from one current lead detector; test was successful;
  • ramp to nominal (4310 A) with 5 A/s ramp rate: NO quench; a FPA was provoked from power converter, with an apparently too high ramp rate which triggered the quench detector.

Proposal for the activity of tomorrow.

Q4 and D2 will be tested from CCC.

Works towards powering the main circuits in the arc and the 600 A correctors (before reaching 1.9 K):

  • plastic bags on 600 A leads (action S. Kockorowski);
  • IST of QPS for which 1.9 K is not required (action R. Denz);
Once 1.9 K are reached:
  • ElQA for the three 13 kA circuits and some selected 600 A correctors (list to be provided by HCC; action D. Bozzini);
  • IST of QPS at 1.9 K (action R. Denz);
  • OK from PIC, PC, EE and Current Leads.

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-23 - 12:26

Short Circuit Tests in UA27 - 23rd May

PDF RAT_20070523.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x23x12x21/RAT_20070523.pdf

Microsoft Excel Spreadsheet SOC_UA27_8hrheatrun.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x23x12x21/SOC_UA27_8hrheatrun.xls

  • The Set Of Circuits for the 8hr heat run attached. It includes all the circuits ion UA27 but the 13kA and the ones belonging to the powering subsector XR2

EE - IST    
13kA EE DONE 16-04-07
600A EE DONE 23-04-07
     
Circuits >2kA    
Powering tests PC 4-8kA - 11-05-07
8h run 4-8kA ON GOING 29-05-07
Interlock tests RQD&F ONGOING 02-05-07
Powering tests RQD&F - 24-05-07
Preparation RB.A23 ON GOING 19-04-07
Interlock tests RB.A23 - 29-05-07
Powering tests RB.A23 - 30-05-07
8h run >= 8kA - 04-06-07
     
Circuits < 2kA    
120A powering tests DONE 03-05-07
600ANOEE powering tests ON GOING 08-05-07
600A-EE interlock tests ON GOING 03-05-07
600A-EE powering tests DONE 18-05-07
8 hr run PC < 2kA - 29-05-07
     
24hr heat run test - 07-06-07
     
Polarity Check < 8KA - 30-05-07
AUG- UPS test - 08-06-07

Edit, attach file or add comment


Created by Main.AntonioVergara on 2007-05-23 - 12:01

Powering Tests Sector 78 - Minutes 23/05/2007

PDF RAT_20070524.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x23x11x59/RAT_20070524.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-22 - 22:46

Powering Tests in Sector 7-8: activity of Tuesday 22/05.

The Powering to Nominal procedure was started fro the following circuits:

  PNO.1 PNO.2
RCBCH5.L8B1 O O
RCBCv5.L8B1 O O
RCBYH4.L8B2 O O
RCBYHS4.L8B1 O O

Several problems were faced:

  • cryogenics instabilities which delayed the CRYO START signal:
    • oscillating level on DFBMC
    • high temperature on the first 6 kA in DFBMA (54 K)
  • delay for PNO.2 step for RCBYH4.L8B2, whose origin was:
    • hardware limitation of the PC, which blocked the cycle to 52 A; the PC went on fault and the current drop to zero (intervention of PO to remove the limitation)
    • probably due to the problem just mentioned, a very long PM file was produced which was preventing the cycle from being restarted (manual reset of the PC was needed)
  • RCBYHS4.L8B2 showed a possible NON-CONFORMITY: while preparing this circuit for PNO.1, a high voltage value was detected on a lead (250 mV); a verification of PO showed that this is probably due to an open voltage tap. The circuit has been removed from the SOC.

PROPOSAL FOR WEDNESDAY 23/05:

  • continuation of tests on RCBCH5.L8B1, RCBCv5.L8B1, RCBYH4.L8B2 and RCBYHS4.L8B1;
  • Powering to Nominal procedure for circuits RCBYV4.L8B1, RCBYVS4.L8B1 and RCBYVS4.L8B2;
  • continuation of tests on Q4 and Q5 if released by MPP;
  • continuation of tests on D2.

POSSIBLE SCENARIOS FOR NEXT WEEK (N.B.: PC, PIC, QPS, EE systems and experts have to be ready for next Tuesday!!!)

RB.A78, RQF.A78 (?), RQD.A78 (?) Powered by their own converter DFBA, QPS, EE, PC, PIC Preparation (this week)
ARC @ 1.9K ElQA @ 1.9K RB & RQD
QPS Heater Discharge
PIC1 RB & RQF
Connection DC cable - lead
PC_CONFIG RQ - 350A
PIC 2 RQ - 350A
PC_CONFIG RB - 350A
PIC 2 RB - 350A
P2N RQ - Max 760A
P2N RB - Max 760A
 
Powered by a 120A converter DFBA, QPS, EE, PC, PIC Preparation (this week)
ARC @ 4.5K PIC1 RB & RQD - Without Quench Heaters
Connection DC cable - lead
PC_CONFIG RQ - 100A
PC_CONFIG RB - 100A
 
600A Circuits - Priority 1, 25 Circuits ARC @ 4.5K or 1.9K DFBA, QPS, EE, PC, PIC Preparation (this week)
ElQA @ 4.5K-1.9K first Circuits
QPS Last Conditioning
PC_CONFIG first circuits
ARC @ Superfluid Helium DFBA, QPS, EE, PC, PIC Preparation (this week)
ElQA @ 4.5K-1.9K All Circuits
QPS Last Conditioning
PIC2 First Circuits
PCS First Circuits
PC_Config, PIC2, PCS rest of circuits - Parallel

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-22 - 19:44

Short Circuit Tests - 21st May

PDF RAT_20070521.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x22x19x40/RAT_20070521.pdf

PDF sct_UA27_v005.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x22x19x40/sct_UA27_v005.pdf

  • SCT planning attached

EE - IST    
13kA EE DONE 16-04-07
600A EE DONE 23-04-07
     
Circuits >2kA    
Powering tests PC 4-8kA - 11-05-07
8h run 4-8kA ON GOING 29-05-07
Interlock tests RQD&F - 02-05-07
Powering tests RQD&F - 24-05-07
Preparation RB.A23 ON GOING 19-04-07
Interlock tests RB.A23 - 29-05-07
Powering tests RB.A23 - 30-05-07
8h run >= 8kA - 04-06-07
     
Circuits < 2kA    
120A powering tests DONE 03-05-07
600ANOEE powering tests ON GOING 08-05-07
600A-EE interlock tests ON GOING 03-05-07
600A-EE powering tests DONE 18-05-07
8 hr run PC < 2kA - 29-05-07
     
24hr heat run test - 07-06-07
     
Polarity Check < 8KA - 30-05-07
AUG- UPS test - 08-06-07

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-22 - 18:34

RAT in Point 4 - Minutes of 22nd May 2007

PDF RAT_20070522.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x05x22x18x33/RAT_20070522.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-22 - 18:33

RAT in Point 4 - Minutes of 11th May 2007

PDF RAT_20070511.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x05x22x18x32/RAT_20070511.pdf

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-22 - 17:47

Warm Magnets Commissioning Point 8 - List of activities for this week (21)

Comments:

List of activities agreed at the Warm Magnets Commissioning Point 8 meeting on 23rd May (8:30 in 936/R-035)

  WM8 LHCb
23.05 MORNING Verification signalization in the tunnel.

Consignation all TI8 LHC PC.

Disconnection of the DC cables to the TI8 LHC PC.

WELQA on the MSI magnets
 
   
   
   
   
   
   
23.05 AFTERNOON WELQA on the rest of the TI8 LHC magnets.

Connection of the DC cables to the TI8 LHC PC.

Déconsignation all TI8 PC.
 
24.05 MORNING Verification signalization in the tunnel.

Repeat the WELQA in the circuits where it could not be completed.

Connection of the DC cables to the TI8 LHC PC.

Déconsignation all TI8 PC.

Polarity check TI8 LHC magnets.

From 13:00 to 14 :30

Déconsignation PC compensators L8:

1. RPTN.SR8.RBXWSH.L8
1. RPTN.SR8.RBXWH.L8

Switch on RPTN.SR8.RBXWSH.L8 and verification at the magnet level.

Switch off RPTN.SR8.RBXWSH.L8

Switch on RPTN.SR8.RBXWH.L8 and verification at the magnet level.

Switch off RPTN.SR8.RBXWH.L8







From 13:00 to 14:30
No work around the ANY compensator magnets
24.05 AFTERNOON Start 24h Heat Run (TI8 LHC + COMPENSATORS L8) (I ultimate) No work around the compensator magnets L8
25.05
MORNING
24h Heat Run ongoing

Temperatures measurements at the end 24h Heat Run
No work around the compensator magnets L8
25.05
AFTERNOON
Turn on all TI8 magnets PC, RPTN.SR8.RBXWSH.L8 and RPTN.SR8.RBXWH.L8.

Start 1st Interferences Evaluation

Start solution on non conformity on power converter of MQIF87800 (3e-4 ripple)
No work around the compensator magnets L8

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-22 - 11:42

Powering Tests Sector 7-8 - Minutes 22/05/2007

PDF RAT_20070523.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x22x11x37/RAT_20070523.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-22 - 00:23

Powering tests of Sector 7-8: activity of Monday, 21

Morning activity. The following tests have been performed on Q5 (MANUAL execution):

  • verification of p-p voltage vs ramp rate (problems with high ramp rates had been discovered last week): ramping up to 200A with 5 A/s, to 400A 7 A/s and 600A with 9 A/s;
  • HCA PLI3.1a: ramp to 2500 A with 5 A/s ramp rate;
  • HCA PLI3.1b, unbalanced cycle: starting from 120 A, one converter ramped to 2500 A at 5 A/s, the other ramped to 1375 at 2.5 A/s;
  • HCA PLI3.1c: skipped after agreement between experts from the different systems;
  • HCA PLI3.4: no icing confirmed by Amalia;
  • HCA PNO.5: ramp to nominal (4310 A); during the ramp, and at the current value of 3941 A, a quench occurred (analysis on going).
Afternoon activity. The following tests have been performed on Q4 (SEQUENCER execution):
  • HCA PLI1.1:
    • first test aborted because the FGC server didn't autorized lhcop to access the power supplies (intervention of S.Page);
    • test stopped, this time because of a cmw time out exception;
    • test failed again because the software current limit have not been set correctly in the FGC;
    • test OK;
  • HCA PLI1.4: FPA at 250 A;
  • HCA PLI1.7: ice absence verified;
  • HCA PLI2.1: current to interm1 (600 A); current reached without problems, but during ramping down the sequencer blocked because of an attemp to reduce the needed time to stand-by
  • HCA PLI2.1: repeated, but during the ramp, around 300 A, a quench occurred (investigations on going).

Proposal for tomorrow:

  • MPP will analyse the data collected on Q4 and Q5, with the feedback from QPS experts;
  • the test activity will focus on 120 A correctors.

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-21 - 10:38

Powering Tests Sector 7-8 - Minutes 21/05/2007

PDF RAT_20070522.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x21x10x36/RAT_20070522.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-20 - 22:20

Powering Tests - 1000A in Q5 -

Done on Friday 18.05.2007

9:30 Powering Procedure for RQ5B1(B2).L8 started
11:30 PLI1.1-4-7 I_inj: 350A
PLI2.1-3-4 I_int_1: 1000A
procedures performed manually and NOT using the sequencer
13:00 Boil-off test on DFBMA-C

Remarks on the work of the last week (14-18 May)

The current was ramped to 1000 A, and the magnet was quenched by triggering the firing of the quench heaters.
The different teams performed a preliminary analysis of the data that had been taken, and some observations were reported:
The DECOUPLING BETWEEN THE TWO CONVERTERS WAS EXCELLENT. By applying a modulation on one power converter of 10A at a DC current of 1000A it was measured that the coupling is negligible (ppm level). The modulation was seen by the quench detectors, but was below the threshold. For future tests, it was suggested to include the current modulation into the procedures, if required (Action: D.Nisbet).
The balancing of the QUENCH DETECTORS was very good. The THRESHOLD was set to 20mV, and the detectors did not trigger. However, it was noted that the PEAK AMPLITUDE OF THE BALANCED SIGNAL WAS SLIGHTLY ABOVE 20mV, BUT FOR A TIME LESS THAN THE VALIDATION TIME OF THE QUENCH DETECTORS OF 10ms. This effect depends on the speed of the ramp, and decreased between the first ramp at 13A/s AND THE SECOND RAMP AT 10A/s.
FOR THE FUTURE COMMISSIONING OF Q5 A RAMP OF 5A/S WILL BE APPLIED, in order to reduce the noise level for the quench detectors. The modified value should be integrated into the sequence. In the second phase the sensitivity of the quench detectors as a function of ramp speed will be further investigated. It might be required to increase the quench detector threshold from 20mV to a higher value. The objective is to ramp the magnet with the nominal rate of about 10A/s.
At the first glance it appears that all required data after the quench were recorded and written into the PM system.
The heater discharge worked as expected. It was correctly recorded, but due to one point with the wrong time the start of graph was wrong. This will be corrected (Action: R.Denz and A.Rijllart).
THE DATA FROM THE CURRENT LEADS (VOLTAGE ETC.) LOOKED OK. It was noticed that the voltage was high during the trigger of the heaters. This is due to the fact that the quench heaters are triggered using a test signal, and that the effect does not indicate an increased voltage across the leads.
AFTER TRIGGERING THE QUENCH, THE HELIUM LEVEL IN THE Q5 MAGNET CRYOSTAT DECREASED IN A SHORT TIME (~MIN) TO ZERO, and increased in about the same time to the nominal value. No change of temperature and pressure were observed. The valve opened after the quench. Since similar changes of the helium level were recorded at other occasions without quenching magnets, it is questions if this effect is real. This needs follow-up, since such drop of helium level would trigger a power abort (Action: ACR). OTHERWISE NO CONSEQUENCES OF THE QUENCH WERE OBSEREVED IN THE CRYOGENIC SYSTEM.
Since the procedures were executed without the sequencer, that data that was recorded needs to be put into MTF manually (Action: HCC)

Program for Monday 21.05.2007

Continue commissioning of Q5, manually without the sequencer.
Start commissioning of the 120A circuits using the sequencer. This can be done in parallel to Q5 commissioning, although there might be some interference.
When Q5 is commissioned, Q4 and D2 will follow. Before starting to commission D2,y the QPS team needs to intervene in order to correct a problem that was reported after PIC2.

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-05-18 - 17:39

RAT meeting 18/5/2007

PDF RAT_20070518.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x18x17x29/RAT_20070518.pdf

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-18 - 15:01

Warm Magnet Commissining Point 8 - Minutes 10.05.07 and planning

PDF Planning.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x18x14x31/Planning.pdf

PDF WM8minutes-presentation.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x18x14x31/WM8minutes-presentation.pdf

PowerPoint Warm_magnets_commissioning_point_8.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x18x14x31/Warm_magnets_commissioning_point_8.ppt

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-17 - 23:49

Powering Tests - Sector 78 -

Done on Wednesday 16.05.2007

9:30 PIC2 on corrector circuits on DFBMA/C
10:30 PIC2 on corrector circuits on DFBMA/C completed
11:00 Problems on RQ4.L8 solved.
12:00 PCC on RQ4.L8 completed
14:30 Problems on RQ5.L8 solved.
15:20 PCC on RQ5.L8 completed
15:30 PIC2 on D2-Q4-Q5.L8
18:00 PIC2 on RD2-RQ4-RQ5.L8 completed.
18:00 Last test of PIC2 not passed “PIC2.10 Test hardware links”
18:15 Tests end

Schedule for Friday 18.05.2007 in sector 7-8

WHEN CONDITIONS FOR THE ElQA IN THE DFBAs ARE MET:

DFBAO ElQA continuity test
DFBAN ElQA continuity test

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL) IN THE POWERING SUBSECTOR ML8:

Sub sector ML8 ready for Powering to Nominal
D2-Q4-Q5.L8 Powering to nominal up to PLI2 (<1000A) – Manually
Corrector circuit 120A DFBMC-A Powering to nominal –Via sequencer

WHEN CONDITIONS IN ARC ARE MET GO ON WITH THE PROGRAMME FOR 60A CONVERTERS (ORBIT CORRECTORS) AND 120A (Q7 TO Q10 CORRECTORS) AS WELL AS THE ELQAs IN THE DFBAs X: previous day - X : last day - O : Proposal

  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-16 - 17:38

Short Circuit Test UA27 - 16th May

PDF RAT_20070516.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x16x16x53/RAT_20070516.pdf

PDF sct_UA27_v004.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x16x16x53/sct_UA27_v004.pdf

Comments: Note that next 18th May, the intervention of the geometers in the UA27 will go on along the day and a WiFi intervention will take place in the morning. During this intervention a lost of the signal is foreseen during about 15 minutes.

     
13kA EE DONE 16-04-07
600A EE DONE 23-04-07
     
Circuits >2kA    
Powering tests PC 4-8kA ON GOING 09-05-07
8h run 4-8kA - 24-05-07
Interlock tests RQD&F - 22-05-07
Powering tests RQD&F - 23-05-07
Preparation RB.A23 ON GOING 19-04-07
Interlock tests RB.A23 - 29-05-07
Powering tests RB.A23 - 30-05-07
8h run >= 8kA - 04-06-07
Pic interface test - 16-06-07
Circuits < 2kA    
120A powering tests DONE 03-05-07
600ANOEE powering tests ON GOING 08-05-07
600A-EE interlock tests - 22-05-07
600A-EE powering tests DONE 08-05-07
8 hr run PC < 2kA - 24-05-07
     
24hr heat run test - 07-06-07
Polarity test < 8KA - 25-05-07
Polarity Check > 8KA - 05-06-07
AUG- UPS test - 08-06-07

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-16 - 17:02

Warm Magnets Commissioning at point 5 - Minutes 11-05-2007

PDF Planning.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2007x05x16x16x40/Planning.pdf

PDF RAT_planning_20070511_WM5.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2007x05x16x16x40/RAT_planning_20070511_WM5.pdf

Comments:

-- Main.MariaPazCasasLino - 16 May 2007

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-16 - 15:50

Powering Tests Sector 7-8 - Minutes 16/05/2007

PDF RAT_20070516.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x16x15x44/RAT_20070516.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-16 - 12:25

Powering Tests: 16 May 2007

Program for 16.05.2007 - afternoon

14:00 PCC for Q5
15:00 PIC2 for D2,Q4,Q5: Ivan and Boris from the CCC, the rest in the FCC If things go smoothly
16:00 Start P2N, preferably Q5, final decision around 15:00. In the sequence a quench at 1000A will be provoked (if we reach this far)

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-15 - 18:59

Powering test - Sector 78 - Tuesday 15th

Done on Tuesday 15.05.2007

9:30 Test start on D2 – QH on.
9:40 Up to ~90A apparently without QPS protection – Actually It was not the case.
15.24 QH enabled for RD2.L8, QPS interlock problem on D2 understood. Only the displayed ref. voltage was a sum of the voltage picked up from the two apertures instead of a difference. At the level of protection the signal was right.
15:34 PCC start on D2 - up to 200A at 16:00
17:30 Loss of Cryo-Start and Cryo-Maintain due to a hardware intervention. Physical cryo conditions were stable. Signals recovered after few minutes.
16:23 PCC for D2 finished successfully
17:40 HQ of RQ4.L8 on – Reiner changed
17:50 Still problems on Q4 – Picked voltage exceed 20mV when PC tries to start up – QPS trips the PC.
   
19:00 Tests end

Schedule for Wednesday 16.05.2007 in sector 7-8

WHEN CONDITIONS FOR THE ElQA IN THE DFBAs ARE MET:

DFBAO ElQA continuity test
DFBAN ElQA continuity test

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL) IN THE POWERING SUBSECTOR ML8:

RQ4.L8 AT-MEL General re-check
AB-PO - Special test with a 60A power converter – charatizarion
DFBMC-A PIC2 on the 120A corrector circuits.

PC CONFIGURATION IN RQ4

DFBMA RQ4.L8
RD2.L8

Balisage with flashing lights HCC
RQ4.L8
RD2.L8
Warn the QPS team that the main circuit will be powered HCC
RQ4.L8
RD2.L8
Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ4.L8
RD2.L8
Unlock and switch on the corresponding power converters AB/PO
RQ4.L8
RD2.L8
PC configuration
UP TO 100 A
AB/PO
PC CONFIGURATION IN RQ5
DFBMC RQ5.L8

Balisage with flashing lights HCC
RQ5.L8 Warn the QPS team that the main circuit will be powered HCC
RQ5.L8 Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ5.L8 Unlock and switch on the corresponding power converters AB/PO
RQ5.L8 PC configuration
UP TO 100 A
AB/PO
WHEN CONDITIONS IN ARC ARE MET GO ON WITH THE PROGRAMME FOR 60A CONVERTERS (ORBIT CORRECTORS) AND 120A (Q7 TO Q10 CORRECTORS) AS WELL AS THE ELQAs IN THE DFBAs X: previous day - X : last day - O : Proposal
  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-15 - 16:13

Powering Tests Sector 7-8 - Minutes 15/05/2007

PDF RAT_20070515.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x15x16x12/RAT_20070515.pdf

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-15 - 14:28

Short Circuit Test UA27 - 14th May

PDF sct_UA27_v003.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x15x14x19/sct_UA27_v003.pdf

PDF RAT_20070514.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x15x14x19/RAT_20070514.pdf

Comments: Note that next 16th May, there will not be powering tests and the PIC interface tests will take place as well as an intervention of the geometers in the UA27 along the day.

Note as well that next 18th a WiFi intervention will take place in the morning. During this intervention a lost of the signal is foreseen during about 15 minutes.

     
13kA EE DONE 16-04-07
600A EE DONE 23-04-07
     
Circuits >2kA    
Powering tests PC 4-8kA ON GOING 09-05-07
8h run 4-8kA - 24-05-07
Interlock tests RQD&F - 16-05-07
Powering tests RQD&F - 23-05-07
Preparation RB.A23 ON GOING 19-04-07
Interlock tests RB.A23 - 29-05-07
Powering tests RB.A23 - 30-05-07
8h run >= 8kA - 04-06-07
Pic interface test - 16-06-07
Circuits < 2kA    
120A powering tests DONE 03-05-07
600ANOEE powering tests ON GOING 08-05-07
600A-EE interlock tests - 16-05-07
600A-EE powering tests DONE 08-05-07
8 hr run PC < 2kA - 24-05-07
     
24hr heat run test - 07-06-07
Polarity test < 8KA - 25-05-07
Polarity Check > 8KA - 05-06-07
AUG- UPS test - 08-06-07

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-15 - 09:54

Beam Dump System Installation and Commissioning, RAT 08 May 2007

PDF RAT_20070508.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2007x05x15x09x53/RAT_20070508.pdf

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-14 - 20:42

Powering test - Monday 14.05 -

Done on Monday 14.05.2007

9:30 PCC on 120A correctors of DFBMC/A
10:30 Deconsignation of Q5-Q4 PC
11:00 Problems with Q5 – Water faults – TS/CV came at 13:30 – they cleaned the filters – Problem recovered.
14.30 AB-PO started to tests D2 with a 60A PC connected to the magnet (circuit definition)– DC cable disconnected from PC side.
14:42 Logging not working. Need to disactivate and re-activate the SOC from SOC editor to restore it. Monitoring problem RCBCH5 and RCBCV5 stopped transmitting data to the monitoring system. Expects contacted. System restarted. Problem disappeared. Systected 'middleware' problem in the network - problem is known and happens once a week.
14.50 Create and activate new SOC named PCC_IPDQ140507 with circuits RQ4.L8 RQ5.L8 and RD2.L8.
15.30 Deconsignation of QH for Q5-Q4-D2
16.30 PCC tests for SOC PCC_070510 finished (corrector circuits DFBMA/C)
16:40 Testing communication between cryo-maintain signal and PIC, DFBMA,DFBMC
16.45 End of the RD2.L8 charaterization
17.30 RQ4.L8B1 and RQ4.L8B2 ready for PCC
17.30 QPS detecting voltages of around ~40-50mV from the magnet protection and trip the PC (not current leads). These trips occur immediately after attempting to ramp the power converters. Note that the QPS trip threshold was reduced from 100mV to 20mV.
18.30 Same problem on Q5.
19:00 Tests end

Schedule for Tuesday 15.05.2007 in sector 7-8

WHEN CONDITIONS FOR THE ElQA IN THE DFBAs ARE MET:

DFBAO ElQA continuity test
DFBAN ElQA continuity test

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL) IN THE POWERING SUBSECTOR ML8:

PC CONFIGURATION IN RQ4, RD2 AND RQ5

DFBMA RQ4.L8
RD2.L8

Balisage with flashing lights HCC
RQ4.L8
RD2.L8
Warn the QPS team that the main circuit will be powered HCC
RQ4.L8
RD2.L8
Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ4.L8
RD2.L8
Unlock and switch on the corresponding power converters AB/PO
RQ4.L8
RD2.L8
PC configuration
UP TO 100 A
AB/PO

DFBMC RQ5.L8

Balisage with flashing lights HCC
RQ5.L8 Warn the QPS team that the main circuit will be powered HCC
RQ5.L8 Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ5.L8 Unlock and switch on the corresponding power converters AB/PO
RQ5.L8 PC configuration
UP TO 100 A
AB/PO
WHEN CONDITIONS IN ARC ARE MET GO ON WITH THE PROGRAMME FOR 60A CONVERTERS (ORBIT CORRECTORS) AND 120A (Q7 TO Q10 CORRECTORS) AS WELL AS THE ELQAs IN THE DFBAs X: previous day - X : last day - O : Proposal
  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

**

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-14 - 10:54

Powering Tests Sector 7-8. Minutes 14/05/2007

PDF RAT_20070514.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x14x10x53/RAT_20070514.pdf

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-11 - 16:26

Schedule for Monday 14.05.07 in sector 78

Schedule for Monday 14.05.2007 in sector 7-8

WHEN CONDITIONS FOR THE ElQA IN THE DFBAs ARE MET:

DFBAO ElQA continuity test
DFBAN ElQA continuity test

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL) IN THE POWERING SUBSECTOR ML8:

DFBMA RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
Balisage with flashing lights HCC
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
PC Configuration (test of PC and magnet at minimum current)
UP TO 10 A
AB/PO

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Balisage with flashing lights HCC
RCBCH5.L8B1
RCBCV5.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
PC Configuration (test of PC and magnet at minimum current)
UP TO 10 A
AB/PO

PC CONFIGURATION IN RQ4, RD2 AND RQ5

DFBMA RQ4.L8
RD2.L8

Balisage with flashing lights HCC
RQ4.L8
RD2.L8
Warn the QPS team that the main circuit will be powered HCC
RQ4.L8
RD2.L8
Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ4.L8
RD2.L8
Unlock and switch on the corresponding power converters AB/PO
RQ4.L8
RD2.L8
PC configuration
UP TO 100 A
AB/PO

DFBMC RQ5.L8

Balisage with flashing lights HCC
RQ5.L8 Warn the QPS team that the main circuit will be powered HCC
RQ5.L8 Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ5.L8 Unlock and switch on the corresponding power converters AB/PO
RQ5.L8 PC configuration
UP TO 100 A
AB/PO
WHEN CONDITIONS IN ARC ARE MET GO ON WITH THE PROGRAMME FOR 60A CONVERTERS (ORBIT CORRECTORS) AND 120A (Q7 TO Q10 CORRECTORS) AS WELL AS THE ELQAs IN THE DFBAs X: previous day - X : last day - O : Proposal
  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

Done on Thursday 10.05.2007

9:30 Swith the Power Converters on: RPMBA.UA83.RQT12.L8B1 and RPMBA.UA83.RQT12.L8B2
Beginning of the validation of the sequencer for the procedure of Q2.
Test on the communication between the PIC and the CRYO signals (CRYO start and CRYO maintain).
10:00 Deconsignation of the PCs of the corrector circuits of Q4 and Q5.
CRYO control in the FCR is not available, except a viewer.
A voltage tap in the circuit RCBYHS4.L8B2 has been detected, the problem will be investigated.
11:30 Waiting for solving the voltage problem.
Consignation of the PCs of the circuits: RPHGA.UA83.RQ9.L8B2 RPHGA.UA83.RQ9.L8B1 RPHGA.UA83.RQ10.L8B2 RPHGA.UA83.RQ10.L8B1 in order to allow AT/MEL to install new heaters.
14:00 Consignation of all the circuits of the converters of the Q4 in order to allow Vincent Chareyre to check the problem of the voltage drop in the circuit RCBYHS4.L8B2.
15:00 Vincent Chareyre has fixed the problem.
15:20 start with the PC configuration of the circuit RCBCH5.L8B1.
16:00 16:10 CRYO-OK lost.
16:20 CRYO-OK found again.
PC configuration of the circuit RCBCH5.L8B1 is ongoing.
17:00 Start with the PC configuration of the RCBCV5.L8B2.
17:10 CRYO-maintain forced LOST and PIC removed all the Power permits. Consequently all the PCs tripped off.
17:30 Cryo Operator reported on the situation: the CRYO OK was lost and the CRYO maintain was going to be lost, consequently they forced the CRYO-MAINTAIN as lost.
18 :00 Switch the Power Converters off: RPMBA.UA83.RQT12.L8B1 and RPMBA.UA83.RQT12.L8B2.

Comments:

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-11 - 11:27

Powering Tests Sector 78 - RAT Minutes - 11/05/2007

PDF RAT_20070511_2.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x11x11x24/RAT_20070511_2.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-10 - 18:50

Schedule for Friday 11.05.2007 in sector 7-8

Schedule for Friday 11.05.2007 in sector 7-8

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL) IN THE POWERING SUBSECTOR ML8:

PC CONFIGURATION IN THE Q4 AND Q5 CORRECTOR CIRCUITS

DFBMA RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
Balisage with flashing lights HCC
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
PC Configuration (test of PC and magnet at minimum current)
UP TO 10 A
AB/PO

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Balisage with flashing lights HCC
RCBCH5.L8B1
RCBCV5.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
PC Configuration (test of PC and magnet at minimum current)
UP TO 10 A
AB/PO

PC CONFIGURATION IN RQ4, RD2 AND RQ5

DFBMA RQ4.L8
RD2.L8

Balisage with flashing lights HCC
RQ4.L8
RD2.L8
Warn the QPS team that the main circuit will be powered HCC
RQ4.L8
RD2.L8
Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ4.L8
RD2.L8
Unlock and switch on the corresponding power converters AB/PO
RQ4.L8
RD2.L8
PC configuration
UP TO 100 A
AB/PO

DFBMC RQ4.L8

Balisage with flashing lights HCC
RQ4.L8 Warn the QPS team that the main circuit will be powered HCC
RQ4.L8 Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ4.L8 Unlock and switch on the corresponding power converters AB/PO
RQ4.L8 PC configuration
UP TO 100 A
AB/PO
WHEN CONDITIONS IN ARC ARE MET GO ON WITH THE PROGRAMME FOR 60A CONVERTERS (ORBIT CORRECTORS) AND 120A (Q7 TO Q10 CORRECTORS) AS WELL AS THE ELQAs IN THE DFBAs

X: previous day - X : last day - O : Proposal

  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

REPORT ON THE ACTIVITIES OF THURSDAY 10TH MAY

09:30 Swith on of the Power Converters: RPMBA.UA83.RQT12.L8B1 and RPMBA.UA83.RQT12.L8B2
Beginning of the validation of the sequencer for the procedure of D2.
Test on the communication between the PIC and the CRYO signals (CRYO start and CRYO maintain).
10:00 Deconsignation of the PCs of the corrector circuits of Q4 and Q5.
CRYO control in the FCR is not available, except a viewer.
A voltage tap problem in the circuit RCBYHS4.L8B2 has been detected, the problem will be investigated.
11:30 Waiting for solving the voltage problem.
Consignation of the PCs of the circuits: RPHGA.UA83.RQ9.L8B2 RPHGA.UA83.RQ9.L8B1 RPHGA.UA83.RQ10.L8B2 RPHGA.UA83.RQ10.L8B1 in order to allow AT/MEL to install new heaters.
14:00 Consignation of all the circuits of the converters of the Q4 in order to allow Vincent Chareyre to check the problem of the voltage drop in the circuit RCBYHS4.L8B2.
15:00 Vincent Chareyre has fixed the problem.
15:20 start with the PC configuration of the circuit RCBCH5.L8B1.
16:00 16:10 CRYO-OK operation lost.
16:20 CRYO-OK operation found again.
PC configuration of the circuit RCBCH5.L8B1 is ongoing.
17:00 Start with the PC configuration of the RCBCV5.L8B2.
17:10 CRYO-maintain forced LOST and PIC removed all the Power permits. Consequently all the PCs tripped off.
17:30 Cryo Operator reported on the situation: the CRYO OK was lost and the CRYO maintain was going to be lost, consequently they forced the CRYO-MAINTAIN as lost.
18 :00 Switch off of the Power Converters: RPMBA.UA83.RQT12.L8B1 and RPMBA.UA83.RQT12.L8B2.

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-10 - 18:26

Powering Test Sector 78 - RAT Meeting - 10/05/2007

PDF RAT_20070510.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x10x18x17/RAT_20070510.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-10 - 08:37

Schedule for Thursday 10.05.2007 in sector 7-8

Schedule for Thursday 10.05.2007 in sector 7-8

FINAL CHECK ON CIRCUITS RPMBB.UA83.RCD.A78B1 AND RPMBB.UA83.RCD.A78B2 TO MAKE THEM AVAILABLE FOR SEQUENCER TESTS ON D2 TYPE PROCEDURE.

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL) IN THE POWERING SUBSECTOR ML8:

PC CONFIGURATION IN THE Q4 AND Q5 CORRECTOR CIRCUITS

DFBMA RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
Balisage with flashing lights HCC
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
PC Configuration (test of PC and magnet at minimum current)
UP TO 10 A
AB/PO

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Balisage with flashing lights HCC
RCBCH5.L8B1
RCBCV5.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
PC Configuration (test of PC and magnet at minimum current)
UP TO 10 A
AB/PO

PC CONFIGURATION IN RQ4, RD2 AND RQ5

DFBMA RQ4.L8
RD2.L8

Balisage with flashing lights HCC
RQ4.L8
RD2.L8
Warn the QPS team that the main circuit will be powered HCC
RQ4.L8
RD2.L8
Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ4.L8
RD2.L8
Unlock and switch on the corresponding power converters AB/PO
RQ4.L8
RD2.L8
PC configuration
UP TO 100 A
AB/PO

DFBMC RQ4.L8

Balisage with flashing lights HCC
RQ4.L8 Warn the QPS team that the main circuit will be powered HCC
RQ4.L8 Unlock the corresponding quench heaters power supplies and charge AT/MEL
RQ4.L8 Unlock and switch on the corresponding power converters AB/PO
RQ4.L8 PC configuration
UP TO 100 A
AB/PO
WHEN CONDITIONS IN ARC ARE MET GO ON WITH THE PROGRAMME FOR 60A CONVERTERS (ORBIT CORRECTORS) AND 120A (Q7 TO Q10 CORRECTORS) AS WELL AS THE ELQAs IN THE DFBAs

X: previous day - X : last day - O : Proposal

  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

Done on Monday 07.05.2007

SEQUENCER TESTS FOR THE PROCEDURE Q4Q5

DFBAO RQ6.L8B1
RQ6.L8B2
Balisage around the chimneys HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. Check of the short-circuit and the insulation. AB/PO
Sequencer tests on powering procedures for 6kA magnets with two power converters ("Q4Q5" procedure) AB/CO
Lock and switch off the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters AB/PO

INSULATION TEST AT THE DFBMA : RESULT OK

CIRCUIT 60A RCBV12.L8B1 INTERVENTION BY AT/MEL: CUPPER COIL PLACED AROUND A CURRENT LEAD

DRY BAGS AND VENTILATORS INTALLATION AT THE DFBMC AND DFBMA BY AT/ACR

“BAD QUALITY OF THE HELIUM LIQUID IN THE LINE CPREVENTED THE POWERING IN ML8 IN THE AFTERNOON”

Done on Tuesday AND Wednesday 08/09.05.2007

GO ON WITH THE VALIDATION OF THE SEQUENCER FOR THE PROCEDURE Q4Q5

DFBAO RQ6.L8B1
RQ6.L8B2
Balisage around the chimneys HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. Check of the short-circuit and the insulation. AB/PO
Sequencer tests on powering procedures for 6kA magnets with two power converters ("Q4Q5" procedure) AB/CO
Lock and switch off the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters AB/PO

INTERVENTION OF THE GEOMETERS AT D2 Q4 (Beginning AT 9:00)

MEASUREMENTS AT THE DFBMs (ML8)

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-09 - 10:48

Short circuit test UA27 - 9th May

PDF RAT_20070509.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x09x10x31/RAT_20070509.pdf

Energy Extraction 600A; Individual Tests are completed. 3 minor problems found (on 16 systems and not 48 as mentionned). Interlock and powering tests completed. Currently installation of temperature- and voltage-probes is on-going. Expected to be completed on May 9. Gert Coelingh

-- Main.BorisBellesia - 09 May 2007

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-05-08 - 17:36

RAT 7/5/2007

PowerPoint RAT2.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x08x17x17/RAT2.ppt
Presentation by A. Ballarino
PDF RAT_20070508.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x08x17x17/RAT_20070508.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-07 - 22:54

Anti-condensation systems tested on DFBMA and DFBMC.

JPEG DSC03896b.jpg:
DSC03896b.jpg
Status of the DFBMA _ Plastic envelopes on all chimneys plus additional heater on top of the 120A-8leads-assembly and air blower in front of it (maximum power).
JPEG DSC03893b.jpg:
DSC03893b.jpg
Status of the DFBMC _ Plastic envelopes on all chimneys plus additional heater on top of both 120A-assemblies and air blower in front of them (minimum power).

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-07 - 22:44

Temperatures, humidity and condensation status in Sector 7-8, Mon 7.

Position T [°C] T_meteo [°C] RH_meteo [%] T_portable [°C] RH_portable [%] Condensation Dew Point [°C] Time
DFBMC   20.6 48.0 20.4 50.8     20.45
Q6_L8 (120A) 9     19.4 52.2 Yes 9.3  
Q8_L8 (120A) 9     19.4 53.1 Yes 9.7 19.45
Q9_L8 (120A) 9.5     19.5 52.7 Yes 9.5  
Q10_L8 (120A) 9     19.0 53.7 Yes 9.4  
Q11_L8 (60A) 9.5     19.0 53.9 Yes 9.4  
Q12_L8 (60A) 9.5 18.6 51.4 18.3 55.5 Low 9.2 20.00
Q13_L8 (60A) 10     18.5 55.0 Yes 9.3  
Q14_L8 (60A) 9.5     18.4 55.4 Yes 9.3  
Q15_L8 (60A) 8.5     18.2 56.0 Yes 9.2 20.15

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-05-07 - 19:34

RAT 7/5/2007

PDF RAT_20070507_last.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x07x19x23/RAT_20070507_last.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-07 - 19:12

Programme for Tuesday 08.05.2007 in Sector 78

Schedule for Thursday 08.05.2007 in sector 7-8

GO ON WITH THE VALIDATION OF THE SEQUENCER FOR THE PROCEDURE Q4Q5

DFBAO RQ6.L8B1
RQ6.L8B2
Switch on the flashing lights around the chimneys of RQ6.L8 circuit in the DFBAO HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. AB/PO
Sequencer tests on powering procedures for 6kA magnets with 2 PCs ("Q4Q5" procedure) AB/CO

INTERVENTION OF THE GEOMETERS AT D2 Q4 (Beginning AT 9:00)

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL), EITHER IN THE DFBMA OR DFBMC, BEGIN WITH THE POWERING TESTS IN ML8

DFBMA RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
RQ4.L8B1

Balisage with flashing lights HCC
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
PC Configuration (test of PC and magnet at minimum current) AB/PO
RQ4.L8 Warn the QPS team that RQ4 will be powered HCC
RQ4.L8 Unlock the quench heaters power supplies of Q5 and charge AT/MEL
RQ4.L8 Unlock and switch on the corresponding power converters AB/PO
RQ4.L8 PC configuration AB/PO
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
RQ4.L8B1
PIC2 HCC
AB/CO
AB/PO
AT/MEL
AT/ACR
RCBYVS4.L8B1
RCBYVS4.L8B2
RCBYHS4.L8B1
RCBYHS4.L8B2
RCBYV4.L8B1
RCBYH4.L8B2
RQ4.L8B1
P2N
(Up to PL1.31B)
HCC
AB/CO
AB/PO
AT/MEL
AT/ACR

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
RQ5.L8
Balisage with flashing lights HCC
RCBCH5.L8B1
RCBCV5.L8B2
Unlock and switch on the corresponding power converters AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
PC Configuration (test of PC and magnet at minimum current) AB/PO
RQ5.L8 Warn the QPS team that RQ5 will be powered HCC
RQ5.L8 Unlock the quench heaters power supplies of Q5 and charge AT/MEL
RQ5.L8 Unlock and switch on the corresponding power converters AB/PO
RQ5.L8 PC configuration AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
RQ5.L8
PIC2 HCC
AB/CO
AB/PO
AT/MEL
AT/ACR
RCBCH5.L8B1
RCBCV5.L8B2
RQ5.L8
P2N
(Up to PL1.31B)
HCC
AB/CO
AB/PO
AT/MEL
AT/ACR

WHEN CONDITIONS IN ARC ARE MET GO ON WITH THE PROGRAMME FOR 60A CONVERTERS (ORBIT CORRECTORS) AND 120A (Q7 TO Q10 CORRECTORS) AS WELL AS THE ELQAs IN THE DFBAs

X: previous day - X : last day - O : Proposal

  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

Done on Monday 07.05.2007

SEQUENCER TESTS FOR THE PROCEDURE Q4Q5

DFBAO RQ6.L8B1
RQ6.L8B2
Balisage around the chimneys HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. Check of the short-circuit and the insulation. AB/PO
Sequencer tests on powering procedures for 6kA magnets with two power converters ("Q4Q5" procedure) AB/CO
Lock and switch off the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters AB/PO

INSULATION TEST AT THE DFBMA : RESULT OK

CIRCUIT 60A RCBV12.L8B1 INTERVENTION BY AT/MEL: COPPER COIL PLACED AROUND A CURRENT LEAD

DRY BAGS AND VENTILATORS INTALLATION AT THE DFBMC AND DFBMA BY AT/ACR

“BAD QUALITY OF THE HELIUM LIQUID IN THE LINE C PREVENTED THE POWERING IN ML8 IN THE AFTERNOON”

Comments:

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-07 - 18:54

Short Circuit Test UA27 - 7th May

PDF RAT_20070507.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x07x18x51/RAT_20070507.pdf

EE - IST    
13kA EE DONE 16-04-07
600A EE DONE 23-04-07
     
Circuits >2kA    
Powering tests PC 4-8kA - 11-05-07
8h run 4-8kA - 29-05-07
Interlock tests RQD&F - 02-05-07
Powering tests RQD&F - 24-05-07
Preparation RB.A23 ON GOING 19-04-07
Interlock tests RB.A23 - 29-05-07
Powering tests RB.A23 - 30-05-07
8h run >= 8kA - 04-06-07
     
Circuits < 2kA    
120A powering tests DONE 03-05-07
600ANOEE powering tests ON GOING 08-05-07
600A-EE interlock tests ON GOING 03-05-07
600A-EE powering tests ON GOING 18-05-07
8 hr run PC < 2kA - 29-05-07
     
24hr heat run test - 07-06-07
     
Polarity Check < 8KA - 30-05-07
AUG- UPS test - 08-06-07

Edit, attach file or add comment


Created by Main.MatteoSolfaroli on 2007-05-07 - 18:38

RAT in Point 4 - Minutes of 4th May 2007

PDF RAT_20070504.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint4x2007x05x07x18x37/RAT_20070504.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-04 - 23:08

Temperatures, humidity and condensation status in Sector 7-8, Fri 4.

Position T [°C]Sorted ascending T_meteo [°C] RH_meteo [%] T_portable [°C] RH_portable [%] Condensation Dew Point [°C] Air Speed [m/s] Time
Q9_L8 (120A) 10     19.3 54.4 No 10.0 1.1-1.4  
Q13_L8 (60A) 10     18.4 55.2 No 9.2    
Q6_L8 (120A) 8.5     19.7 53.1 Low 9.9    
Q15_L8 (60A) 8.5     18.2 55.9 Yes 9.2   19.15
Q8_L8 (120A) 9     19.8 53.2 Low 10.0    
Q10_L8 (120A) 9     19.0 54.2 Low 9.5   18.50
Q11_L8 (60A) 9     19.0 54.6 No 9.5    
Q12_L8 (60A) 9 18.5 52 18.6 54.5 Yes 9.2 1.0-1.3  
Q14_L8 (60A) 9     18.5 55.1 Yes 9.3    
DFBMC   20.4 48.1         0.3-0.8 17.30
    20.4 48.3 20.1 51.1       19.30

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-04 - 19:27

Short Circuit Test in UA27- Schedule:

EE - IST    
13kA EE DONE 16-04-07
600A EE DONE 23-04-07
     
Circuits >2kA    
Powering tests PC 4-8kA - 11-05-07
8h run 4-8kA - 29-05-07
Interlock tests RQD&F - 02-05-07
Powering tests RQD&F - 24-05-07
Preparation RB.A23 ON GOING 19-04-07
Interlock tests RB.A23 - 29-05-07
Powering tests RB.A23 - 30-05-07
8h run >= 8kA - 04-06-07
     
Circuits < 2kA    
120A powering tests ON GOING 03-05-07
600ANOEE powering tests - 08-05-07
600A-EE interlock tests ON GOING 03-05-07
600A-EE powering tests - 18-05-07
8 hr run PC < 2kA - 29-05-07
     
24hr heat run test - 07-06-07
     
Polarity Check < 8KA - 30-05-07
AUG- UPS test - 08-06-07

Edit, attach file or add comment


Created by Main.BorisBellesia on 2007-05-04 - 18:55

SCT in UJ33

PowerPoint GlobalAnalysis24hHR_UJ33.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint3x2007x05x04x18x27/GlobalAnalysis24hHR_UJ33.ppt

  • During the 24hr heat run major problems were found (for a more detailed description see the attached file):
1 - Over-heating of the systems in UJ33 (central part) – TS-EL will re-routing the cables inside the UJ33.

2 - Over- heating of the cables in the tunnel side RZ33 – a vertical cable tray will be positioned parallel to the TS-CV cooling water supports. The tray will host one of the three cable layers that are now placed on the same tray.

  • The re-routing cable works will last approximately up to the third week of June. The integration/cost studies are on going and the possible co-activities interfering with the preparation for the cooldown of the A34 are being taken into account.
A meeting to up-date the "status of the art" is foreseen for Monday 14th. Details will be given.

  • During the inspection of the UJ33 P. Proudlock commented:

1 - ODH system must be foreseen in UJ33 - the DFBLC inside the UJ makes it indispensable

2 - Rectangular ducts (for the geometers) must be closed during the heat-run

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-04 - 18:35

Programme for Monday 07.05.2007

Programme for Monday 07.05.2007

DFBAO RQ6.L8B1
RQ6.L8B2
Switch on the flashing lights around the chimneys of RQ6.L8 circuit in the DFBAO HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. AB/PO
Sequencer tests on powering procedures for 6kA magnets with 2 PCs ("Q4Q5" procedure) AB/CO

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL)

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Balisage with flashing lights HCC
PC Configuration (test of PC and magnet at minimum current) AB/PO
RQ5.L8 Unlock of the quench heater power supplies (DQHDS) of Q5 and charge. AT/MEL
PC Configuration AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
RQ5.L8
PIC2 HCC
AB/CO
AB/PO
AT/MEL
AT/ACR
P2N HCC
AB/CO
AB/PO
AT/MEL
AT/ACR

WHEN CONDITIONS IN ARC ARE MET
go on with the programme for the 60A converters (orbit correctors) and 120A (Q7 to Q10 correctors) as well as ElQAs in DFBAs

X: previous day - X : last day - O : Proposal

  PC_Config PNO.1 PNO.2 PNO.3  
RCBH11_L8B1 X X X X  
RCBV11_L8B2 X X X X  
RCBH12_L8B2 X X X O  
RCBV12_L8B1 X X X O  
RCBH13_L8B1 X X X O  
RCBV13_L8B2 X X X O  
RCBH14_L8B2 X X X O  
RCBV14_L8B1 X X X O  
RCBH15_L8B1 X X O O  
RCBV15_L8B2 X X O O  
RCBH16_L8B2 X X O O  
RCBV16_L8B1 X X O O  
RCBH17_L8B1 -       PC_Config/crowbar tests to be repeated
RCBV17_L8B2 -       PC_Config/crowbar tests to be repeated
RCBH18_L8B2 -       PC_Config/crowbar tests to be repeated
RCBV18_L8B1 -       PC_Config/crowbar tests to be repeated
RCBH23_L8B1 O        
RCBV23_L8B2 O        
RCBH24_L8B2 O        
RCBV24_L8B1 O        
RCBH25_L8B1          
RCBV25_L8B2          
RCBH27_L8B1          
RCBV27_L8B2          
RCBH28_L8B2          
RCBV28_L8B1          
RCBH29_L8B1          
RCBV29_L8B2          
RCBH31_L8B1          
RCBV31_L8B2          
RCBH32_L8B2          
RCBV32_L8B1          

  PC_Config PIC2 PNO.1 PNO.2 PNO.3 PNO.4  
RCBCH9_L8B1 -           PC_Config/crowbar tests and TFA missing
RCBCV9_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCH10_L8B2 -           PC_Config/crowbar tests and TFA missing
RCBCV10_L8B1 -           PC_Config/crowbar tests and TFA missing

Done on Friday 04.05.2007

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Connection of the DC cables to the power converters for these circuits (before 9:00) AB/PO
Verification of the connections at the DFBMC side AT/MEL
DFBAO RQ6.L8B1
RQ6.L8B2
Switch on the flashing lights around the chimneys of RQ6.L8 circuit in the DFBAO HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. AB/PO
Sequencer tests on powering procedures for 6kA magnets with 2 PCs ("Q4Q5" procedure) AB/CO

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-05-04 - 16:35

RAT meeting 4/5/2007

PowerPoint Current_lead_operating_conditions_20070504.ppt:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x04x16x33/Current_lead_operating_conditions_20070504.ppt
Presentation by A. Ballarino - conditions for powering
PDF RAT_20070504.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x04x16x33/RAT_20070504.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2007-05-04 - 16:03

Cooldown Sector 78

Despite the long series of faults (network, cooling water, pumping group) experienced at the end of last week, the cryo team managed to bring down the average temperature of the magnets to 3 K during the week-end. The switch over to the new cryoplant took place on Monday but the connection of the new pumping group was delayed until Wednesday morning. Stabilization of temperatures and pressures as well as fine adjustments of the plants took place most of Wednesday until Thursday at 11.30 when the pumping group was set at 15 mbar. Unfortunately, at 11.30 an attempt to restart the UPS system on the surface caused a power cut which stopped all the PLCs. The cryo team is still recovering from that failure and we expect to be at nominal cryogenic condition in the arc tonight Friday May 4th.

Following a top management meeting which took place on Monday, a series of measures towards the clearing of the difficulties with the commissioning of the DFBs were taken: two different ways to avoid condensation on the chimneys are now in place: blowers on 6kA leads + heaters on the 120A leads on the DFBMC and dry-air bags around the chimneys of the DFBMA.

Although the powering front was very quiet during the week, significant progress was achieved with the preparation of the sequencer for the independently powered quadrupoles and the 120A correctors: we now have the sequencer tuned for the 60 A and 120 A correctors as well as the independently powered quadrupoles and D2.

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-03 - 23:46

Anti-Condensation Systems tested on DFBMA and DFBMC

JPEG DSC03835b.jpg:
DSC03835b.jpg
DFBMA: Additional heater on the 120A-8leads-assembly and plastic envelope with forced air on the 120A-4leads-assembly
JPEG DSC03830b.jpg:
DSC03830b.jpg
DFBMC: Detail of the air blowers in front of the 6 kA chimneys/leads
JPEG DSC03828.JPG:
DSC03828.JPG
DFBMC: Additional heaters on the 120A chimneys and air blowers in front of the 6kA chimneys/leads
JPEG DSC03840b.jpg:
DSC03840b.jpg
DFBMA: Plastic envelopes with forced air on the 6 kA

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-03 - 22:55

Temperatures, relative humidity and condensation status in Sector 7-8, Thu 3.

Position T [°C] T_meteo [°C] RH_meteo [%] Condensation Comments
DFBMC   20.7 38.8    
Q6_L8 (120A) 9     Yes  
Q8_L8 (60A) 9.5     No  
Q9_L8 (60A) 13     No  
Q10_L8 (60A) 10     Yes  
Q11_L8 (60A) 9.5     No  
Q12_L8 (60A) 9 18.5 41.4 Yes  
Q13_L8 (60A) 10     No  
Q14_L8 (60A) 9     No  
Q15_L8 (60A) 9     Yes  
Q20_L8 (60A) 9     No  
Q25_L8 (60A) 10     No  
Q30_L8 (60A) 10     No  
Q30_R7 (60A) 9.5     Yes  
Q25_R7 (60A) 10.5     No  
Q20_R7 (60A) 10     No  
Q15_R7 (60A) 9     Yes  
Q14_R7 (60A) 9     Yes  
Q13_R7 (60A) 9.5     Yes  
Q12_R7 (60A) 9.5 17.3 49 Yes  
Q11_R7 (60A) 9     Yes  
Q10_R7 (60A) 9     Yes  
Q9_R7 (60A) 9.5     Yes  
Q8_R7 (60A) 9     Yes  
Q7_R7 (60A) 7     Yes  
DFBAN   18 45.9    

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-03 - 18:03

RAT meeting 3/5/2007

Programme for Friday 04.05.2007

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Connection of the DC cables to the power converters for these circuits (before 9:00) AB/PO
Verification of the connections at the DFBMC side AT/MEL
DFBAO RQ6.L8B1
RQ6.L8B2
Switch on the flashing lights around the chimneys of RQ6.L8 circuit in the DFBAO HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. AB/PO
Sequencer tests on powering procedures for 6kA magnets with 2 PCs ("Q4Q5" procedure) AB/CO

WHEN CONDITIONS FOR POWERING ARE MET (CRYO, MPP, MEL)

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Balisage with flashing lights HCC
PC Configuration (test of PC and magnet at minimum current) AB/PO
RQ5.L8 Unlock of the quench heater power supplies (DQHDS) of Q5 and charge. AT/MEL
PC Configuration AB/PO
RCBCH5.L8B1
RCBCV5.L8B2
RQ5.L8
PIC2 HCC
AB/CO
AB/PO
AT/MEL
AT/ACR
P2N HCC
AB/CO
AB/PO
AT/MEL
AT/ACR

Done on Thursday 03.05.2007

DFBMC RCBCH5.L8B1
RCBCV5.L8B2
Connection of the DC cables to the DFBMC leads of the two 120A circuits which passed the ElQA tests AT/MEL
DFBAO RQ6.L8B1
RQ6.L8B2
Balisage around the chimneys HCC
Unlock and switch on the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters. Check of the short-circuit and the insulation. AB/PO
Sequencer tests on powering procedures for 6kA magnets with two power converters ("Q4Q5" procedure) AB/CO
Lock and switch off the RPHGB.UA83.RQ6.L8B1 and RPHGB.UA83.RQ6.L8B2 converters AB/PO


Unlock and switch on the RPLB.UA83.RCBCVS5.L8B1 converter for sequencer Tests 80-120 A AB/PO

Sequencer tests 80-120A AB/CO AB/CO

DFBAO Installation of dry-air bags on the DFBMA chimneys AT/ACR
DFBMC Installation of two heaters in two 120 A chimneys and of the air blowers for the 6kA leads AT/ACR

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-03 - 13:41

Short Circuit Test in UA27-02May

PDF RAT_20070502.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x03x13x36/RAT_20070502.pdf

PDF sct_UA27_v002.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint2x2007x05x03x13x36/sct_UA27_v002.pdf

Comments:

Edit, attach file or add comment


Created by Main.MikeKoratzinos on 2007-05-03 - 12:59

RAT meeting 3/5/2007

PDF RAT_20070503.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x03x12x58/RAT_20070503.pdf

Comments:

Edit, attach file or add comment


Created by Main.MariaPazCasasLino on 2007-05-03 - 12:52

Planning of the day 03.05.2007

PDF Planning_of_the_day_03.05.2007.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x05x03x12x49/Planning_of_the_day_03.05.2007.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2007-05-03 - 02:01

Temperatures, relative humidity and condensation status in Sector 7-8, Wed 2.

Position T [°C] T_meteo [°C] RH_meteo [%] Condensation Comments
DFBMA          
120A-4 14     No Setpoint 40K - additional heaters
120A-8 13     No Setpoint 50K - additional heaters
6kA 8 (on top of G11)   Yes No condensation on chimneys - Setpoint 50K
DFBMC   20.5 44.2    
6kA 8 (on G11)   No Setpoint 70K
120A-8 11     No Setpoint 50K - additional heaters not working
120A-4 11     No Setpoint 40K - additional heaters not working
DFBAO_LCM          
Q6_L8 (120A) 8     Yes  
DFBAO_HCM          
Q9_L8 (60A) 12     No  
Q10_L8 (60A) 11     No  
Q11_L8 (60A) 10     No  
Q12_L8 (60A) 9 18.3 49.5 Yes  
Q13_L8 (60A) 11     No  
Q14_L8 (60A) 9     Yes  
Q15_L8 (60A) 10     Yes  
Q20_L8 (60A) 9     Yes  
Q25_L8 (60A) 11     Yes  
Q30_L8 (60A) 11     Yes  
Q30_R7 (60A) 10     -  
Q25_R7 (60A) 11     -  
Q20_R7 (60A) 10     No  
Q15_R7 (60A) 9     Yes  
Q14_R7 (60A) 9     Yes  
Q13_R7 (60A) 15?     No  
Q12_R7 (60A) 10 17.3 53.4 Yes  
Q11_R7 (60A) 9     Yes  
Q10_R7 (60A) 9     Yes  
Q9_R7 (60A) 10     No  
Q8_R7 (60A) 9     Yes  
Q7_R7 (60A) 9     Yes  
DFBAN   18 49.9 Yes  
600A 10     Yes Setpoint 150K
13kA 10     Yes Setpoint 110K
DFBMH       No  
600A 11     No Setpoint 120K

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-05-02 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login